OpenStudy (anonymous):

1+1=2 so why is it so complicated to do the rest need help LOL! urgently The functions f(x) and g(x) are defined below. f(x) = 3x - 6 g(x) = -e^x - 5 Using a table of values, determine the solution to the equation f(x) = g(x). A. x = 2 B. x = 5 C. x = 0 D. x = -6

OpenStudy (anonymous):

f(x)=g(x) \[3x-6=-e ^{x}-5\] \[3x-6+5=-e ^{x},3x-1=-e ^{x}\] \[1-3x=e ^{x}=1+\frac{ x }{ 1! }+\frac{ x ^{2} }{2! }+\frac{ x ^{3} }{ 3! }+...\] \[0=x \left( 4+\frac{ x ^{2} }{2! } +\frac{ x ^{3} }{3! }+...\right)\] hence one solution is x=0

OpenStudy (anonymous):

thank u for explaining this problem do u think youre able to help me with somemore

OpenStudy (anonymous):

I will try.